# Evolutionary Hysteresis: Cycling about in a Rugged Landscape

### v1

@@ VII Data Availability
Original data generated in this study is accessible from the Zenodo repository: 10.5281/zenodo.15784542. Public data used in this study are referenced in the article.
@@ VIII Code Availability
The code developed for simulation and analysis is available in the Zenodo repository: 10.5281/zenodo.15784542. Additional simulation and analysis scripts are also available from the following Github repositories: LukePiszkin/Wright-Fisher-Simulation, /Protein-Evolution-Hysteresis, and /Lake_Mendota_hysteresis. All original code generated for results presented in the study are included in the article, supplementary information, or Zenodo repository.

### v2

@@ VII Data Availability
Original data generated in this study is accessible from the Zenodo repository: 10.5281/zenodo.18001835. Public data used in this study are referenced in the article.
@@ VIII Code Availability
The code developed for simulation and analysis is available in the Zenodo repository: 10.5281/zenodo.18001835. Additional simulation and analysis scripts are also available from the following Github repositories: LukePiszkin/Wright-Fisher-Simulation, /Protein-Evolution-Hysteresis, and /Lake_Mendota_hysteresis. All original code generated for results presented in the study are included in the article, supplementary information, or Zenodo repository.